OpenStudy (anonymous):

Given v = (5, -2) and u = (6, -4), find 3v and 2u The answers are: (2,-6) (-9,17) (27,-14) (6,13)

OpenStudy (zzr0ck3r):

3v just means multiply each entry of v by 3 so \[3v=3\left(\begin{matrix}5 \\ -2\end{matrix}\right)\]=\[\left(\begin{matrix}15 \\ -6\end{matrix}\right)\] I dont see this in your answer list so you must have posted the question wrong??? maybe 3v+2u = [15,-6]+[12,-8]=[27,-14]

OpenStudy (anonymous):

I get that it means multiply. All the other questions I was able to get the correct answer, but this one doesn't seem to add up.

OpenStudy (anonymous):

Nevermind, I see what I did...Thanks

OpenStudy (zzr0ck3r):

np